You want to install a hardwood floor in a rectangular room. The floor has a perimeter of 76 feet and a length of 25 feet.
(a) What is the width of the floor?

Answer by Maths68(1474)

Length = l= 25 feet
Width = w=?
Perimeter = p = 76 feet
==============================
Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle =2 length +2 width
P = 2l +2 w
Put the values of p,l and w in above equation
76=2(25)+w
76=50+2w
76-50=2w
26=2w
2w=26
Divide by 2 both sides of above equation
2w/2=26/2
w=13
Width of the floor = 13 feet